6.

0.50 points Problem 10-10 Calculating Real Returns and Risk Premiums Youve observed the following returns on Yasmin Corporations stock over the past five years: 16 percent 19 percent, 13 percent, and 10 percent. Suppose the average inflation rate over this periad was 2.2 percent and the average T-bill rale over the period was 5.3 percent. alroun od to 2 decimal pl a. What wras the average real return on Yasmins stock? (Do not round intenmediate calculations. Enter your answer as a percent rounded to 2 decimal places (e.g, 32.16).] Average real return$6 b. What was the average nominal risk premium on Yasmins stock? (Do not round intermediate calculations. Enter your answer as a percent rounded to 1 decimal place (e.g., 32.1).) erage nominal risk
Answer #1

To calculate the average real return, we need to calculate the average nominal returns over the period and also use average inflation rate in the fischer equation.

According to Fischer equation

(1+ nominal rate of return) = (1+real rate of return) (I+inflation rate)

Average Nominal rate = (16%-5%+19%+13%+10%) / 5
      = 10.60%
Given Average Inflation rate = 2.2%

Putting it in fischer equation

(1+10.60%) = (1+ real rate of return) * (1+2.2%)
(1+0.1060) = (1 + real rate of return) * (1+ 0.022)
1.1060/1.022 = 1+ real rate of return
1.082191 - 1 = real rate of return
real rate of return = 0.082191
      = 8.22%

Average real rate of return = 8.22%

b)Average Nominal risk premium = Average nominal rate of return - Nominal risk free rate
      = 10.60% - 5.3%
      = 5.30%

Average nominal risk premium = 5.30%